# Control-Oriented Deep Space Communications For Unmanned Space Exploration

### v2

@@ III Control-Oriented Optimization of the Mother-Daughter Downlink
Restricted by payload limitations, the spacecraft usually carries lightweight BSs, which are less powerful than ground-based facilities. The restricted transmit power of the onboard BS, coupled with the cycle-time constraint leads to the limited cycle rate of the mother-daughter downlink. This influences the accuracy of transmit commands and affects the control performance. To alleviate the bottleneck effect of mother-daughter downlink in the $\mathbf{SC^{3}}$ loop, we propose a control-oriented communication scheme. This scheme optimizes the block length and transmit power of the mother-daughter to minimize the sum LQR cost of the $K$ $\mathbf{SC^{3}}$ loops,
@@
where $R(p_{k})\triangleq R(n_{k}^{*},p_{k})$ and $L(p_{k})\triangleq L(n_{k}^{*},p_{k})$. To solve this problem, we first analyze their monotonicity and convexity-concavity. For the cycle rate $R(p_{k})$, we analyze the property of the achievable rate expression, i.e., $r(p_{k})\triangleq r(n_{k}^{*},p_{k})$, as given in (8), which holds the same property as $R(p_{k})$. For ease of expression, we rewrite $r(p_{k})$ as a function of the signal-to-noise ratio (SNR), denoted as $\gamma_{k}\triangleq\frac{g_{k}p_{k}}{\sigma^{2}}$,
@@ IV Comparisons With Two Communication-Oriented Schemes
2:  Calculate $\mathbf{S}_{k}$, $\mathbf{M}_{k}$, $\mathbf{N}_{k}$, and $\mathbf{\Sigma}_{k}$ according to (II)-(21);
@@ IV-A Communication-Oriented Power Allocation
$\sum_{k=1}^{K}p_{k}\leqslant P_{\max}.$ (37c)
where (36b) and (37b) are derived from the stable condition (30) and $p_{k}^{i-1}$ is the solution obtained in the $(i-1)$ th iteration. The algorithm is assured to converge for the monotonicity and boundedness of generated solutions. We summarize the corresponding algorithm in Algorithm 2.
@@
2:  Calculate $\mathbf{S}_{k}$, $\mathbf{M}_{k}$, $\mathbf{N}_{k}$, and $\mathbf{\Sigma}_{k}$ according to (II)-(21);
@@ IV-B Approximate Closed-Form Solutions
In this subsection, we derive the approximate closed-form solutions for three discussed schemes. We assume that all the $\mathbf{SC}^{3}$ loops have the same maximum allowable transmission time, i.e., $T_{k}=T,\forall k$. According to (24), the optimal block length is calculated by $n=\lfloor BT\rfloor$. In addition, on the condition that the received SNR is greater than 5 dB, i.e., $10\log_{10}(\frac{g_{k}p_{k}}{\sigma^{2}})\geqslant 5$ dB, we can approximate the cycle rate, i.e., $R(p_{k})$, as follows [31]
@@
Furthermore, when the system is in the assure-to-be-stable region, where the cycle rate is much larger than the intrinsic entropy, i.e., $R(p_{k})\gg\log_{2}|\det\mathbf{A}_{k}|$, the rate-cost function has the following approximation:
@@
$=\frac{m|\det\mathbf{N}_{k}\mathbf{M}_{k}|^{\frac{1}{m}}|\text{det}\mathbf{A}_{k}|^{\frac{2}{m}}e^{\frac{2\sqrt{n}}{m}Q^{-1}(\epsilon_{k})}}{(1+\frac{gp_{k}}{\sigma^{2}})^{\frac{2n}{m}}}$
@@
By testing the Slater condition, it is easy to find that the strong duality holds for (P3). Solving its dual problem yields the same optimal solution. The dual problem is given by
@@
Moving forward, we derive the closed-form solutions for two communication-oriented schemes. Using the high-SNR approximation in (38), (PA) and (PB) become two convex problems. As for (PA), if without the stable condition (32b), the optimal solution is given by the classical Water-Filling method [32, Ch 9.4]
@@ V Simulation Results and Discussion
In the simulation, we assume that there are $K=4$ $\mathbf{SC^{3}}$ loops. For sensing-related parameters, we set $\mathbf{C}_{k}=\mathbf{I}_{m}$ and $\mathbf{\Sigma}_{\mathbf{w}_{k}}=\mathbf{0}_{m}$. For control-related parameters, we set $m=100$, $\mathbf{Q}_{k}=\mathbf{I}_{m}$, $\mathbf{R}_{k}=\mathbf{0}_{m}$, and $\mathbf{\Sigma}_{\mathbf{v}_{k}}=0.01*\mathbf{I_{m}}$ [22]. The intrinsic entropy is given by $\log_{2}|\det\mathbf{A}_{k}|=\{48,36,24,12\}$ for $k=1\sim 4$. The control cycle time is $T_{k}^{0}=50$ ms and the time used for sensing, computing and control is $40$ ms. In this way, the time allowed for communication is 10 ms, which accords to the closed-loop control requirement in space exploration [33]. As for communication-related parameters, each $\mathbf{SC^{3}}$ loop is allocated with a narrow frequency band of $B=15$ kHz. The carrier frequency is $f_{c}=2$ GHz, the variance of the channel noise is $\sigma^{2}=-110$ dBm [22], and the transmission error probability is $\epsilon_{k}=10^{-6}$. The maximal antenna gain is $G_{\max}=23$ dBi and $\theta_{\text{3dB}}=30^{\circ}$. Given that there are no high buildings on the explored planet, we use the rural scenario to model the near-planet environment. The channel-related parameters are given by $P_{Los}=91.9\%$, $SF_{1}=1.14$ dB, $SF_{2}=8.78$ dB and $CL=18.42$ dB [27]. According to the height of the Mars orbit, we set the height of the mother spacecraft to 3000 km and four probes are located $[1000,2000,3000,4000]$ km away from the horizontal projection point of the spacecraft. In the present figures, the dotted line represents the infinite LQR cost and the system is unstable in such cases.
@@
In Fig. 4, we compare the proposed scheme with two control-oriented schemes [21][22]. The scheme in [21] defined a new objective named the energy-to-control efficiency. We thereby use $\frac{P_{\max}-\sum\limits_{k=1}^{K}{p_{k}}}{\sum\limits_{k=1}^{K}L(n_{k},p_{k})}$ as its objective in the simulation. As for the scheme proposed in [22], it took the LQR cost as the objective, but this scheme used the Shannon capacity to calculate the cycle rate. It can be seen that the scheme proposed in [22] performs the worst when $P_{\max}\leqslant 4.5$ dBw. This is attributed to its unawareness of the rate losses in the FBL regime. Some loops that should have been allocated with more power are under allocated. This leads to the system instability in the low power region. As the power increases, the LQR cost under the scheme in [21] becomes the highest. This is due to its emphasis on energy efficiency. The control performance has to make a compromise to save power.
@@
In Fig. 5, we compare the proposed scheme with three communication-oriented schemes: the max-sum rate scheme, the max-min rate scheme, and the classical equal power allocation. The curves under the approximate closed-form solution of the proposed scheme (44) and the modified max-min rate scheme (PC) are also shown. It can be seen that the control performance using the approximate closed-form solution (44) is nearly the same as that using the accurate solution. The performance under the modified max-min rate scheme shows a small gap from the optimal one. These results confirm their excellent approximations to the optimal solution. In addition, we can also see that the max-min rate scheme outperforms equal power allocation. Both of them further outperform the max-sum rate scheme. This is because of their different power allocation principles in terms of the channel condition. The max-min rate scheme holds a similar allocation principle as the proposed scheme, while the max-sum rate scheme follows an opposite allocation principle.
@@
In Fig. 6, we present the power allocation results of the proposed scheme, the max-sum rate scheme, and the max-min rate scheme. The maximal transmit power is set as $P_{\max}=10$ dBw. It is evident that in both subfigures, the power ratio increases from loops one to four under the proposed scheme and the max-min rate scheme, while it decreases under the max-sum rate scheme. This verifies their power allocation principles in terms of channel conditions. In addition, when the cycle time is $50$ ms, the proposed scheme allocates more power to $\mathbf{SC^{3}}$ loops $1\sim 3$ compared with the max-min rate scheme. This is because we set $\log_{2}|\det\mathbf{A}_{k}|=\{48,36,24,12\}$ for $k=1\sim 4$. This adjustment is made to compensate for the intrinsic entropy difference. In the top subfigure, which shows the case of $T_{k}^{0}=5$ s, the power allocation results of the proposed scheme and the max-min rate scheme are nearly the same. In this situation, the intrinsic entropy difference has little impact on the communication power allocation. This verifies our conclusions drawn in Theorem 3.
@@
Next, we discuss the impacts of the cycle time on the control performance. The results are presented in Fig. 7. It can be seen that the cycle time impacts the control performance in the lower power region. As the power increases, different curves converge to the same minimal value. This is because in the low–power region, the cycle rate is very limited. Different cycle time brings with different cycle rates and different accuracy of the commands, which, consequently, leads to different control performances. While in the high-power region, all loops have sufficient cycle rates to deliver commands accurately. The control performances all go to the same optimum, making the cycle time difference less influential in this region.
@@
In Fig. 8, we present the impacts of the sensing noise on the control performance. Different from the impact of cycle time, the LQR cost under different sensing noise would not converge to the same minimal value as the power increases. This is because inaccurate sensing leads to the control command deviating from the optimal one. This deviation cannot be eliminated regardless of how excellent the communication is. In fact, in the flat region of these curves, the bottleneck of the $\mathbf{SC}^{3}$ loop is sensing but not communication.
@@ Appendix A Proof of Theorem 1
$\log_{2}(\gamma_{0}+1)$ $-\log_{2}(e)\eta\sqrt{1-\frac{1}{(\gamma+1_{0})^{2}}}=0$ (58)
@@
{$&r(\gamma)\ \text{is convex-concave,}\quad\eta\geqslant\frac{(\hat{\gamma}+1)\ln(\hat{\gamma}+1)}{\sqrt{\hat{\gamma}^{2}+2\hat{\gamma}}},\\ &r(\gamma)\ \text{is concave,}\qquad\qquad\quad\eta>\frac{(\hat{\gamma}+1)\ln(\hat{\gamma}+1)}{\sqrt{\hat{\gamma}^{2}+2\hat{\gamma}}}.\end{aligned}\blacksquare\right.$
@@ Appendix B Proof of Theorem 2
For simplicity, we rewrite the cycle rate and the rate-cost function as the functions of SNR, i.e., $R(\gamma)$ and $L(\gamma)$. In addition, a new function is introduced as follows
@@
$\triangledown^{2}\Omega(\gamma)$ $*\displaystyle\leqslant n\log_{2}e\bigg{(}-\frac{1}{(\gamma+1)^{2}}+\frac{\log_{2}(\gamma+1)}{\log_{2}e}*$ (80)

### v3

@@ III Control-Oriented Optimization of the Mother-Daughter Downlink
Next, we propose our control-oriented optimization scheme. Compared with those schemes addressed for general IoT scenarios, our formulated problem address the extreme conditions in the space environment. Specifically, given the limited communication resources carried by the space robots, we directly take the control performance as the optimization objective. Additionally, concerning the varying environment in space, the data transmission within the $\mathbf{SC}^{3}$ loop is characterized in the FBL regime. Based on these two special concerns, our formulated problem thus takes the downlink block length and transmit power as the optimization variables to minimize the sum LQR cost of $K$ $\mathbf{SC^{3}}$ loops:
@@
where $R(p_{k})\triangleq R(n_{k}^{*},p_{k})$ and $L(p_{k})\triangleq L(n_{k}^{*},p_{k})$. To solve this problem, we first analyze their monotonicity and convexity-concavity. For the cycle rate $R(p_{k})$, we analyze the property of the achievable rate expression, i.e., $r(p_{k})\triangleq r(n_{k}^{*},p_{k})$, as given in (8), which holds the same property as $R(p_{k})$. For ease of expression, we rewrite $r(p_{k})$ as a function of the SNR, denoted as $\gamma_{k}\triangleq\frac{g_{k}p_{k}}{\sigma^{2}}$,
@@ IV Comparisons With Two Communication-Oriented Schemes
2:  Calculate $\mathbf{S}_{k}$, $\mathbf{M}_{k}$, $\mathbf{N}_{k}$, and $\mathbf{\Sigma}_{k}$ according to (19)-(21);
@@ IV-A Communication-Oriented Power Allocation
$\sum_{k=1}^{K}p_{k}\leqslant P_{\max},$ (37c)
where (36b) and (37b) are derived from the stable condition (30) and $p_{k}^{i-1}$ is the solution obtained in the $(i-1)$ th iteration. The algorithm is assured to converge because of the monotonicity and boundedness of generated solutions. We summarize the corresponding algorithm in Algorithm 2.
@@
2:  Calculate $\mathbf{S}_{k}$, $\mathbf{M}_{k}$, $\mathbf{N}_{k}$, and $\mathbf{\Sigma}_{k}$ according to (19)-(21);
@@ IV-B Approximate Closed-Form Solutions
In this subsection, we derive the approximate closed-form solutions for three discussed schemes. We assume that all the $\mathbf{SC}^{3}$ loops have the same maximum allowable transmission time, i.e., $T_{k}=T,\forall k$. According to (24), the optimal block length is calculated by $n=\lfloor BT\rfloor$. In addition, when the received SNR is greater than $10$ dB, i.e., $10\log_{10}(\frac{g_{k}p_{k}}{\sigma^{2}})\geqslant 10$ dB, the term $\sqrt{V(p_{k})}$ in (10) is very close to 1, i.e., $\sqrt{V(p_{k})}|_{SNR\geqslant 10dB}\geqslant 0.9959$. Under this condition, the cycle rate expression can be approximated as follows [41]:
@@
Furthermore, if the system further works in the assure-to-be-stable region, that the cycle rate is much larger than the intrinsic entropy, i.e., $R(p_{k})\gg\log_{2}|\det\mathbf{A}_{k}|$, we can omit the term of minus one in the numerator in the rate-cost function and obtain the following approximation:
@@
By combining (38) and (39), we can get the following approximation of the rate-cost expression:
$L(p_{k})$ $\approx\frac{m|\det\mathbf{N}_{k}\mathbf{M}_{k}|^{\frac{1}{m}}|\text{det}\mathbf{A}_{k}|^{\frac{2}{m}}e^{\frac{2\sqrt{n}}{m}Q^{-1}(\epsilon_{k})}}{(1+\frac{gp_{k}}{\sigma^{2}})^{\frac{2n}{m}}}$ (40)
@@
Based on above approximation processes, we can see that (40) closely aligns to the original expression when the system is in the assure-to-be-stable region. However, this approximation would lose its efficiency when the working point is out of this region. Since the following introduced closed-form solution is derived from the approximation of the rate-cost function, it works well in the assure-to-be-stable region where the accuracy of these approximations are maintained. Correspondingly, the precision of the closed-form solution diminishes in the region approaching the critical stability, where the approximations are no longer accurate. By testing the Slater condition, it is easy to find that the strong duality holds for (P3). Solving its dual problem yields the same optimal solution. The dual problem is given by
@@
Moving forward, we derive the closed-form solutions for two communication-oriented schemes. Using the high-SNR approximation in (38), (PA) and (PB) become two convex problems. As for (PA), if without the stable condition (32b), the optimal solution is given by the classical Water-Filling method [42, Ch. 9.4]
@@ V Simulation Results and Discussion
In the simulation, we assume that there are $K=4$ $\mathbf{SC^{3}}$ loops. For sensing-related parameters, we set $\mathbf{C}_{k}=\mathbf{I}_{m}$ and $\mathbf{\Sigma}_{\mathbf{w}_{k}}=\mathbf{0}_{m}$. For control-related parameters, we set $m=100$, $\mathbf{Q}_{k}=\mathbf{I}_{m}$, $\mathbf{R}_{k}=\mathbf{0}_{m}$, and $\mathbf{\Sigma}_{\mathbf{v}_{k}}=0.01*\mathbf{I_{m}}$ [29]. By setting $\mathbf{R}_{k}=\mathbf{0}_{m}$, the LQR cost quantifies state deviations, which directly reflects the control effectiveness of the $\mathbf{SC}^{3}$ loop. This accords to our focus of the control effectiveness rather than the control cost. In the case of $\mathbf{R}_{k}\neq\mathbf{0}_{m}$, the LQR cost changes accordingly per (19). Corresponding simulation results will give the same conclusions as in the case of $\mathbf{R}_{k}=\mathbf{0}_{m}$, and we do not further discuss this case. The intrinsic entropy is given by $\log_{2}|\det\mathbf{A}_{k}|=\{48,36,24,12\}$ for $k=1\sim 4$. The control cycle time is $T_{k}^{0}=50$ ms and the time used for sensing, computing and control is $40$ ms. In this way, the time allowed for communication is 10 ms, which accords to the closed-loop control requirement in space exploration [43]. As for communication-related parameters, each $\mathbf{SC^{3}}$ loop is allocated with a narrow frequency band of $B=15$ kHz. The carrier frequency is $f_{c}=2$ GHz, the variance of the channel noise is $\sigma^{2}=-110$ dBm [29], and the transmission error probability is $\epsilon_{k}=10^{-6}$. The maximal antenna gain is $G_{\max}=23$ dBi and $\theta_{\text{3dB}}=30^{\circ}$. Given that there are no high buildings on the explored planet, we use the rural scenario to model the near-planet environment. The channel-related parameters are given by $P_{Los}=91.9\%$, $SF_{1}=1.14$ dB, $SF_{2}=8.78$ dB and $CL=18.42$ dB [16]. According to the height of the Mars orbit, we set the height of the mother spacecraft to 3000 km and four probes are located $[1000,2000,3000,4000]$ km away from the horizontal projection point of the spacecraft. In the present figures, the dotted line represents the infinite LQR cost and the system is unstable in such cases.
Figure 4: Comparisons of the LQR cost under the accurate solution and the proposed closed-form solution.
In Fig. 4, we discuss the applicability of the proposed approximate closed-form solution (45). In this simulation, the minimal transmit power to stabilize the control system is 3.49 dBw. It can be seen that when $P_{\text{sum}}<8\text{dBw}$, the LQR cost under the closed-form solution is obvious higher than that under the optimal solution. This is because the approximations of the cycle rate (38) and the rate-cost function (39) deviate from their original expressions. In contrast, when $P_{\text{sum}}\geqslant 8\text{dBw}$, the LQR cost under the closed-form solution is nearly the same as that under the optimal solution. This verifies the applicability of the closed-form solution in the assure-to-be-stable region.
@@
In Fig. 5, we compare the proposed scheme with two control-aware schemes for general IoT scenarios [28][29]. In [28], the authors paid particular attention to the low-latency communication but they did not prioritize the task performance in the optimization design, using a hybrid metric named the energy-to-control efficiency. We thereby use $\frac{P_{\max}-\sum\limits_{k=1}^{K}{p_{k}}}{\sum\limits_{k=1}^{K}L(n_{k},p_{k})}$ as its objective for a fair comparison. In contrast, the scheme proposed in [29] directly optimized control LQR cost, but it overlooked the cycle-time constraint, applying the Shannon capacity to calculate the cycle rate. Since these two schemes were either not completely control oriented or time sensitive, the control performance of corresponding “mother-daughter system” is inferior compared with the proposed scheme, making them less effective in the space setting. It can be seen that the scheme proposed in [29] performs the worst when $P_{\max}\leqslant 4.5$ dBw. This is attributed to its unawareness of the rate losses in the FBL regime. Some loops that should have been allocated with more power are under allocated. This leads to the system instability in the low-power region. As the power increases, the LQR cost under the scheme in [28] becomes the highest. This is due to its emphasis on the energy efficiency. The control performance has to make a compromise to save power. These results validate the effectiveness of the proposed scheme in the delay-sensitive and resource-constrained space setting.
@@
In Fig. 6, we compare the proposed scheme with three communication-oriented schemes: the max-sum rate scheme, the max-min rate scheme, and the classical equal power allocation. The curves under the approximate closed-form solution of the proposed scheme (45) and the modified max-min rate scheme (PC) are also shown. The performance under the modified max-min rate scheme shows a small gap from the optimal one. These results confirm their excellent approximations to the optimal solution. In addition, we can also see that the max-min rate scheme outperforms equal power allocation. Both of them further outperform the max-sum rate scheme. This is because of their different power allocation principles in terms of the channel condition. The max-min rate scheme holds a similar allocation principle as the proposed scheme, while the max-sum rate scheme follows an opposite allocation principle.
@@
In Fig. 7, we present the power allocation results of the proposed scheme, the max-sum rate scheme, and the max-min rate scheme. The maximal transmit power is set as $P_{\max}=10$ dBw. It is evident that in both subfigures, the power ratio increases from loops one to four under the proposed scheme and the max-min rate scheme, while it decreases under the max-sum rate scheme. This verifies their power allocation principles in terms of channel conditions, recalling $d_{k}=[1000,2000,3000,4000]\ \text{km}$. In addition, when the cycle time is $50$ ms, the proposed scheme allocates more power to $\mathbf{SC^{3}}$ loops $1\sim 3$ compared with the max-min rate scheme. This is because we set $\log_{2}|\det\mathbf{A}_{k}|=\{48,36,24,12\}$ for $k=1\sim 4$. This adjustment is made to compensate for the intrinsic entropy difference. In the top subfigure, which shows the case of $T_{k}^{0}=5$ s, the power allocation results of the proposed scheme and the max-min rate scheme are nearly the same. In this situation, the intrinsic entropy difference has little impact on the communication power allocation. This verifies our conclusions drawn in Theorem 3.
@@
Next, we discuss the impacts of the cycle time on the control performance. The results are presented in Fig. 8. It can be seen that the cycle time impacts the control performance in the low-power region. As the power increases, different curves converge to the same minimal value. This is because in the low–power region, the cycle rate is very limited. Different cycle time brings with different cycle rates and different accuracy of the commands, which, consequently, leads to different control performances. While in the high-power region, all loops have sufficient cycle rates to deliver commands accurately. The control performances all go to the same optimum, making the cycle time difference less influential in this region.
@@
In Fig. 9, we present the impacts of the sensing noise on the control performance. In this simulation, we set the sensing noise covariance as $\sigma^{2}_{w}=[0,1e^{-3},2e^{-3}]$. Different from the impact of cycle time, the LQR costs under different sensing noises would not converge to the same minimal value as the power increases. This is because the inaccurate sensing would lead the control command deviating from the optimal one. This deviation cannot be eliminated regardless of how excellent the communication is. This can also be explained by the rate-cost function (22). The function’s last term, i.e., $\text{tr}(\mathbf{\Sigma}_{k}\mathbf{A}_{k}\mathbf{M}_{k}\mathbf{A}_{k})$, is determined by the sensing noise and control parameters and is irrelevant of the communication quality. The larger value of $\sigma_{w}^{2}$ leads to the larger value of $\text{tr}(\mathbf{\Sigma}_{k}\mathbf{A}_{k}\mathbf{M}_{k}\mathbf{A}_{k})$, which can not be minimized by good communication. In fact, in the flat region of these curves, the value of the limit LQR cost is determined by the last two terms of the rate-cost function. In this situation, the bottleneck of the $\mathbf{SC}^{3}$ loop is sensing but not communication.
@@ Appendix A Proof of Theorem 1
$\log_{2}(\gamma_{0}+1)$ $-\log_{2}(e)\eta\sqrt{1-\frac{1}{(\gamma_{0}+1)^{2}}}=0$ (59)
@@
{$&r(\gamma)\ \text{is convex-concave,}\quad\eta\leqslant\frac{(\hat{\gamma}+1)\ln(\hat{\gamma}+1)}{\sqrt{\hat{\gamma}^{2}+2\hat{\gamma}}},\\ &r(\gamma)\ \text{is concave,}\qquad\qquad\quad\eta>\frac{(\hat{\gamma}+1)\ln(\hat{\gamma}+1)}{\sqrt{\hat{\gamma}^{2}+2\hat{\gamma}}}.\end{aligned}\blacksquare\right.$
@@ Appendix B Proof of Theorem 2
For simplicity, we rewrite the cycle rate and the rate-cost function as the functions of the SNR, i.e., $R(\gamma)$ and $L(\gamma)$. In addition, a new function is introduced as follows
@@
$\triangledown^{2}\Omega(\gamma)$ $\leqslant n\log_{2}e\bigg{(}-\frac{1}{(\gamma+1)^{2}}+\frac{\log_{2}(\gamma+1)}{\log_{2}e}*$ (81)